OpenStudy (anonymous):

how do you do this problem 11-3 2/8

OpenStudy (anonymous):

Make it an improper fraction. Put 11 over 1 then find a common denominator subtract and simplify

OpenStudy (anonymous):

10 8/8-3 2/8=7 6/8 = 7 3/4

OpenStudy (anonymous):

either method works!

OpenStudy (anonymous):

11 - 3 2/8 = 11 - 26/8 since 3 2/8 resolves to 26/8 Now the problem becomes 11 - 26/8 using 8 as the common denominator, this becomes (88-26)/8 = 62/8 = 7 6/8 or 7 3/4